OpenStudy (anonymous):

what are the solutions to the equation |x - 3 | = 14

OpenStudy (kva1992):

x-3=-14 x=-11 x-3=14 x=17

OpenStudy (anonymous):

x= 17 x=-11

OpenStudy (anonymous):

What are the solutions to the equation |2x 3| + 4 = 17?

OpenStudy (anonymous):

its supposed to be 2x - 3

OpenStudy (kva1992):

when you ask for solutions your asking for the value of x

OpenStudy (kva1992):

so lets say 2x-3 was your answer its not a solution yet you still have to solve for x

OpenStudy (kva1992):

and it would be x= 3/2